TIP: Click on subject to list as thread! ANSI
echo: sb-nasa_news
to: All
from: Hugh S. Gregory
date: 2003-03-19 23:19:00
subject: 3\06 Pt-1 ISS On Orbit Status 06-03-2003

This Echo is READ ONLY !   NO Un-Authorized Messages Please!
 ~~~~~~~~~~~~~~~~~~~~~~~~   ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
3\06 ISS On-Orbit Status 06-03-2003
Part 1 of 2

ISS On-Orbit Status 3/6/03

All ISS systems continue to function nominally, except as noted 
previously or below. Today Expedition 6 has been in charge of the 
Station for 100 days.

After wake-up, the crewmates were congratulated on yesterday's 
successful recovery of the station-wide OpsLAN (see below).

After power-up of the SSRMS (space station remote manipulator system) 
by MCC-H at 9:00am EST, CDR Ken Bowersox and FE-2/SO Don Pettit 
performed the second part of this week's Robotics operations, viz., 
the planned survey of the P1 truss segment, followed by minor 
troubleshooting on the Lab DCP (display and control panel) 
Pause/Proceed switch. [Targets were two RBVM QD (radiator beam valve 
module quick disconnect) thermal covers and the port SVS (space vision 
system) target array. The surveys are to ensure that the protective 
booties are properly in place, lest higher temperatures over longer 
periods degrade the ammonia QD seals underneath. On ¾, Sox and Don had 
finished the S1 survey early enough to perform the SSRMS walkoff and 
base change to the Lab PDGF (power and data grapple fixture) for 
today. The results of that S1 survey are still being assessed, but 
preliminary indications are that two RBVMs may require additional 
inspection.]

FE-1 Nikolai Budarin meanwhile continued de-installing and removing
no-longer-needed SUD motion control system components in the FGB 
"Zarya" module to make room for stowage, this time two electronics 
blocks (M34-20-1 and 2). He then took digital photographs of the 
exposed area for later downlink via OCA.

Bowersox and Pettit completed another test activity of the HRF US 
(Human Research Facility Ultrasound) program, a Space Medicine 
investigation using an ultrasound/Doppler system to obtain images of 
internal organs of the human body in zero-G. [After Dr. Peggy Whitson 
had conducted an operational checkout of the system for medical 
contingencies on 9/13/02, including a cardiac scanning which 
demonstrated the validity of remote guidance and promising quality of 
downlinked data, today's activity was to validate a specific zero-G 
cardiac scanning protocol based on terrestrial echocardiography 
standards. This activity ensures the availability of cardiac imaging 
for any future medical contingencies. The ultrasound equipment is part 
of HRF Rack 1, launched on 5A.1 in March 2001. Its applications 
include echocardiography, abdominal US (deep organ), vascular US, 
muscle and tendon US, transcranial US and US contrast studies.]

Budarin began the new phase of the Russian BIO-5 Rasteniya-2/Lada-2
("Plants-2") experiment which researches growth and development of 
plants under spaceflight conditions. [Nikolai unstowed experimental 
seeds (acacia leaf type pea) and planted them between wicks in a root 
tray. The hardware was then powered up. Regular daily maintenance of 
the experiment includes monitoring of seedling growth, humidity 
measurements and moistening of the substrate, and photo/video 
recording.]

Don Pettit completed the daily routine maintenance of the SM's SOSh 
life support system, including ASU toilet facilities, and later 
prepared the IMS inventory delta file for downlink.

He and Bowersox filled out their weekly FFQ (food frequency 
questionnaire), a special software log on the MEC (medical equipment 
computer) to track nutritional intake.

At 9:55am EST, ISS attitude control was handed over to the Russian MCS
(motion control system), which then (10:03am-10:18am) maneuvered the 
Station from sun-fixed XPOP (x-axis perpendicular to orbit plane) to 
earth-oriented LVLH TEA (local vertical/local horizontal torque 
equilibrium attitude).  Attitude control was then resumed by U.S. CMG 
(control moment gyro) momentum management. [In the new LVLH attitude, 
U.S. GNC (guidance, navigation & control) systems currently continue 
to be configured for Russian attitude, angular rates, and state 
(position, velocity) sources. This is important since in LVLH any 
aberrant control impulses generated due to the current RGA (rate gyro 
assembly) "staleness" problem of the R3 software would be much larger 
even than in XPOP and could possibly lead to loss of CMG attitude 
control.]

Houston and Moscow have agreed on a schedule for the U.S. GNC patch 
uplink and the subsequent Service Module (SM) 7.01 software 
transition. [MCC-H will receive the patch tomorrow (3/7) and perform 
uplink testing on Monday (3/10). The onboard transition will occur as 
part of the nominally planned RS handover and Progress reboost test 
scheduled for 3/11. TsUP/Moscow will then resume the SM 7.01 
transition on about 3/18.]

Planning continues for the Stage EVA by Bowersox and Pettit on 3/20
(alternately, 3/25). [The 6.5-hr. spacewalk will focus on close 
inspections and reconfigurations on the S0 and S1 truss segments, 
reconfiguration of CMG hardware at the Z1 truss, installation of two 
more SPDs (spool positioning devices) on jumper connectors, 
repositioning RBVM thermal booties, installation of a light fixture on 
a boom stanchion on one of the two S1 CETA (crew equipment translation 
aid) railcarts (an activity not completed during the last Stage EVA on 
1/15), and some get-ahead tasks as time permits.]

Full functionality of the Station OpsLAN (operations local area 
network) between U.S. (USOS) and Russian segment (RS) has been was 
restored. [Based on intermediate screen capture files downlinked by 
Budarin, TsUP specialists were able to determine that the 
configuration file on the SM's SmartSwitch router (SSR, Russian: BRI) 
was incorrect, such that the ports where the SSC (station support 
computer) laptops are connected were not enabled or configured for 
use. A new config file was uplinked and successfully installed in the 
SSR computer by the crew. The SM SSCs were then able to establish a 
connection to the USOS file server PC. As of now, the SSC OpsLAN is in 
a good configuration.]

At 8:55am EST, MCC-H payload specialists conducted a teleconference 
with the crew to discuss U.S. manifest priorities for 5S/Soyuz TMA-1 
return, to obtain crew input in the ongoing downmass planning. [Moscow 
expects total 5S return capacity to be on the order of 30-50 kg. 
Houston was expecting more and is working to confirm the real 
capability, as well as whether there is any standard hardware that 
could be removed from 5S before return (e.g.  winter coats, etc). With 
worst-case capacity, U.S. allocation of downmass is expected to be 
less than 10 kg. Our return priorities have been established as: 
science first, then minimal environmental samples, then other hardware 
necessary to troubleshoot or understand a technical problem, and 
finally crew preference.]

At 12:56pm EST, the crew participated in a 20-min. interactive 
educational interview on NASA TV with students at Glenwood Elementary 
School in Perrysburg, Ohio. [The school regularly televises morning 
announcements that relate to space. Students in all grade levels are 
engaged in space related activities, and the school's Computer Club 
was to videotape the program and interview audience members, with the 
tape then to be distributed to other schools within the district. NASA 
GRC (Glenn Research Center) sent a representative to present student 
assemblies before the live program.]

All crewmembers performed their regular daily physical exercise on 
TVIS (treadmill with vibration isolation and stabilization), RED 
(resistive exercise device), CEVIS (cycle ergometer with vibration 
isolation) and VELO bike with load trainer. [On the VELO ergometer, 
loading control was found to be intermittent during exercise 
yesterday. Budarin removed the control panel cover and inspected the 
cabling but found no apparent cause. After closing the cover, the 
ergometer performed nominally. In preparation for any future 
re-occurrence, Nikolai located spare ergometer hardware in the RS.]

(continued)

---
* Origin: SpaceBase[tm] Vancouver Canada [3 Lines] 604-473-9357 (1:153/719)
SEEN-BY: 633/267 270
@PATH: 153/719 715 7715 140/1 106/2000 633/267

SOURCE: echomail via fidonet.ozzmosis.com

Email questions or comments to sysop@ipingthereforeiam.com
All parts of this website painstakingly hand-crafted in the U.S.A.!
IPTIA BBS/MUD/Terminal/Game Server List, © 2025 IPTIA Consulting™.